# Dissecting the SWE-Bench Leaderboards: Profiling Submitters and Architectures of LLM- and Agent-Based Repair Systems

### v1

@@ 3.1.2. Analysis of Submitter Organization
Submission to one or more leaderboards. The Venn diagram in Figure 3 shows which organizations have submitted to SWE-Bench Lite, SWE-Bench Verified, or both. Nineteen contributors (32%) submitted solutions to both leaderboards, most of them (12) with duplicated submissions (conversely, seven contributors submitted different solutions to both leaderboards). For example, Amazon submitted Amazon Q Developer Agent (v20240430-dev) to both leaderboards. On the other hand, 21 contributors have submitted only to SWE-Bench Lite, and 19 exclusively to SWE-Bench Verified. For instance, IBM has submitted only to SWE-Bench Lite (with two entries), while EPAM has submitted only to SWE-Bench Verified (with five entries).
Top Submitters. Table 4 compiles the top submitters to the two leaderboards. The ranking is led by Princeton University, the authors of SWE-Bench themselves (jimenez2024SWEBenchLLMs, ), with 11 entries in SWE-Bench Lite and 12 in SWE-Bench Verified. Their submissions are related to two approaches: (1) the initial evaluation on SWE-Bench using retrieval-augmented generation (RAG) and LLMs (jimenez2024SWEBenchLLMs, ), (2) the agent-based solutions SWE-Agent (yang2024sweagentagent, ). Moreover, each approach is often submitted in multiple variants. The most common difference across entries from the same submitter is the underlying LLM used. For instance, SWE-Agent has been evaluated with several LLMs, including Claude 3 Opus, Claude 3.5 Sonnet, Claude 3.7 Sonnet w/ Review Heavy, GPT-4, and GPT-4o. Other organizations, such as EPAM, the second most active submitter to SWE-Bench Verified, include version identifiers in their submission names (e.g., v20241029 and v20250219).
From this table, we observe that nearly all types of organizations are represented among the top submitters: academia, with Princeton University and the University of Illinois Urbana-Champaign; large companies, including Alibaba and Amazon (each with 4 entries in SWE-Bench Verified), and Bytedance (3 entries in SWE-Bench Lite); medium-sized companies, such as Anthropic (3 entries in SWE-Bench Verified); small companies, like nFactorial (4 entries in SWE-Bench Verified); and open-source communities, exemplified by Moatless (5 entries in SWE-Bench Lite).
3.1.3. Analysis of % Resolver rate
Figure 2(b) shows the evolution of the % Resolved metric, which is the primary measure reported in both leaderboards. In spite a period among Oct. 2024 and Nov. 2024 in which some solutions uploaded to SWE-Bench Verified were not optimal, the precision is notably higher in SWE-Bench Verified than in SWE-Bench Lite. In particular, the median and maximum of both leaderboards are 44.2% and 68.2% vs. 30.5% and 55.0%, respectively. Looking at Table 3, we see that this increase is consistent in all submitter categories, and especially in submitters of mixed industry-academy profile. The table also shows that solutions from academy submitters perform poorly compared to solutions from industry, with a median close to half value in both leaderboards. This difference may be attributed to the construction of SWE-Bench Verified, which was derived from the original SWE-Bench by manually filtering out instances considered too hard or unsolvable, with the aim of increasing the benchmark’s robustness and reliability (chowdhury2024swebenchverified, ).
Table 2 shows the median and maximum % Resolved achieved by each type of submitter across the two leaderboards. In SWE-Bench Lite, the highest precision (55%) is achieved by a small company, Isoform, with an entry of the same name. Next we find submissions from companies of other sizes; for example, the public traded company company Globant achieved 48.33%. Academia reaches a maximum precision of 40.67% with the approach Agentless-1.5 and a lower median precision (20.7%), primarily due to the modest performance of earlier submissions. In fact, there is no single academy-only solution appearing in the top-ten % Resolved ranking [^14] In position 10th we find Lingxi, whose type remains unknown.. However, things improve when considering collaboration with industry, with Sprinklr AI et al. ranking 7th with 47% and three other collaborations ranking 11th-13th. The above-mentioned academy top-ranked Agentless-1.5 does not show up until the 15th position, i.e. in the lower part of the first quartile. It is also worth mentioning that the highest median values correspond to categories with only one data point: a medium-sized company (Blackbox AI) and another entry (Kodu AI) for which company size could not be determined. Excluding these outliers, small companies achieve the highest overall median precision (37%).
We assess whether there are statistically significant differences in precision across submitter types. We apply the Kruskal–Wallis test, and, when significant differences are detected, we follow up with Dunn’s post-hoc test using Bonferroni correction. In SWE-Bench Lite, the Kruskal–Wallis test reveals a statistically significant difference between at least two groups of submitters (H = 28.2867, p = 0.0016). Dunn’s post-hoc test identifies a significant difference between Academia and Small Companies ($pvalue$ = 0.00073).
In SWE-Bench Verified, the results present a different landscape. A publicly traded company, NVIDIA, achieved the highest precision with its approach Nemotron-CORTEXA, reaching 68.2%. However, other categories also attained competitive results. These include other types of companies (e.g., Anthropic), academia (e.g., Princeton University), and the open-source community (e.g., ModelScope), all surpassing 62% precision. Excluding the outlier open-source solution ModelScope, the highest median precision is observed in medium companies, driven by submissions from Anthropic, which evaluated its state-of-the-art LLMs featuring hybrid reasoning capabilities.
In SWE-Bench Verified, the Kruskal–Wallis test returned a similar result as SWE-Bench Lite (H = 25.5840, p = 0.0024). Nevertheless, Dunn’s post-hoc test found that the precision scores from small and medium companies differ significantly from those of academia ($pvalues$ 0.00883 and 0.04414). This finding corroborates the trend discussed above.
@@ 3.1.3. Analysis of % Resolver rate
Fig 5 shows the entries of the leaderboard arranged according to the date and precision displayed in the leaderboards.
@@
Firstly, we observe that the initial entries from both leaderboards come from academia, in particular from the authors of the benchmark. The benchmark was initially used to evaluate the ability of large language models (LLMs) to repair software issues (jimenez2024SWEBenchLLMs, ), and was later applied to assess LLM-based agents (yang2024sweagentagent, ). These early approaches achieved a maximum precision of around 20%. Then, starting in May–June 2024, both leaderboards began receiving submissions from other organizations beyond academia.
Figure 5(a) shows the evolution of submissions in SWE-Bench Lite. We observe that Small companies have been constantly pushing the results. Notably, Isoform achieved a state-of-the-art precision of 55% on Jan. 2025, a benchmark that has yet to be surpassed by later submissions. Nevertheless, at certain points, large publicly traded companies have achieved state-of-the-art results. For example, Alibaba reached 33% precision with its Lingma Agent (ma2024lingmaswegpt, ) on 2024-06-22, while Globant attained 48.3% with its Code Fixer on 2024-11-27. We also observe that, since June 2024, submissions from academia have rarely come close to matching the precision of those submitted by companies. As shown in the figure, entries resulting from collaborations between academia and industry tend to achieve precision levels between those of small companies and academic institutions. This suggests that such collaborations can enhance the performance of academic submissions.
As shown in Fig. 5(b) from SWE-Bench Verified, Small companies have also consistently driven improvements in results. However, the top submissions come from the industry: Nemotron-CORTEXA with 68.20% on 2025-05-16, followed by Aime-Coder 66.40% on 2025-05-14 submitted by the DevInfra team from ByteDance. Surprisingly, the solution proposed by this team was inspired by the “implementation of some tools and prompts of Vote Ensembler” from the Augment Code approach [^15] Aime Coder submission: https://github.com/SWE-bench/experiments/blob/main/evaluation/verified/20250514_aime_coder/README.md. In addition to offering commercial products (e.g., development plugins), the small company Augment Code has open-sourced the agent used in the SWE-Bench evaluation [^16] Augment Code Agent: https://github.com/augmentcode/augment-swebench-agent/tree/main. Notably, this agent held the highest precision on the leaderboard (65.4%) in March 2025. This open-source strategy not only enhances the transparency of their results, but also contributes to the broader progress of the community, as evidenced by its influence on ByteDance’s submission.
Nevertheless, in 2025, other types of companies also reached state-of-the-art precision, for example, the publicly traded company EPAM, with 62.8%, and the medium-sized AI leader Anthropic, with 63.2%. With a few exceptions, such as Agentless-1.5 (which achieved 50.8%), submissions from academia have generally exhibited significantly lower precision than those from companies. This disparity may be attributed to differences in engineering resources, compute availability and variations in team size and funding. Notably, when academia and industry collaborated, the results tended to approach state-of-the-art levels. For instance, the collaboration between the University of California, Santa Barbara (USA) and Meta led to the development of PatchPilot (li2025patchpilot, ), whose version v1.1 achieved a near state-of-the-art precision of 64.6% on 2025-05-03.
@@
Answer to RQ 1 (Submitters): Entries in the SWE-Bench leaderboard originate from a wide variety of submitters, including (combinations or particularization of) academy, industry and individual developers (who may belong to open source communities). Small companies are dominant, but Academy has almost an equal share of entries in SWE-Bench Lite, although the number of distinct submitters is significantly greater in Academia (with an average of 3 entries per submitter) than in the rest of categories (less than 2 entries per submitter in average). This diversity underscores the benchmark’s accessibility and minimal barriers to participation, demonstrating that it is readily approachable for experimentation by a broad spectrum of contributors. Since the introduction of SWE-Bench Verified, companies —especially small ones—have been driving progress toward state-of-the-art performance.
@@ 3.1.4. Product and Accessibility
PAP Coding Assistant IDE plugin 8 12 33.5 7 14 41.5 10 26 Alibaba Lingma Agent, Amazon Q Developer Agent, AppMap Navie, Augment Agent, Blackbox AI Agent, Bytedance MarsCode Agent, CodeShellAgent, CodeShellTester, Kodu-v1, OpenCSG Starship
[…]
Cloud platform 1 1 37 0 0 0 1 1 Patched.Codes Patchwork
@@
[…]
Total 16 21 36 17 32 47.9 24 53
[…]
Total 15 30 26.5 7 19 30.2 17 49
[…]
@@
[…]
Total 2 3 30.7 3 8 42.9 4 11
[…]
@@
Table 5 presents the number of (#A) approaches and (#E) entries per product availability category, along with their median precision (Resolved). We analyse below the contents of each category.
@@
Publicly Available Product (PAP). It includes publicly accessible commercial tools and accounts for 52 entries across both leaderboards, representing 24 distinct approaches. This category include products with different purpose: Coding Assistant, Development Platforms, Development Assistant. The most common product type is Coding Assistant, comprising 10 products available as IDE plug-ins and linked to half of the 52 entries. These include, for example, Amazon Q Developer, Alibaba Lingma, and AppMap Navie, which are evaluated on SWE-Bench and available as extensions for popular development environments such as IntelliJ JetBrains IDE and Visual Studio Code. The second major product type is Development Platforms, typically delivered as cloud-based services rather than IDE plug-ins. These platforms generally allow developers to specify new features using natural language and integrate directly with code repositories (e.g., GitHub). There are 21 entries linked to these platforms across both leaderboards and include solutions such as OpenHands, Solver, and Factory Code Droid. A third type includes Development Assistant, which help the developer in particular development task. Three of the entries, such as CodeStory Aide (later rebranded as AgentFarm), materialize this assistant as automated bots integrated into GitHub workflows, interacting with developers via pull requests, while the two others as a command-line tool and a cloud platform. Notably, the submissions related to a Development Assistant product achieve the highest median % Resolved across both leaderboards.
@@
The Non-Commercial Solutions (NCS) category accounts for 49 entries across both leaderboards, corresponding to 17 distinct approaches. The dominant product type is Issue Resolution, comprising 12 approaches such as Agentless, SWE-agent, HyperAgent, and AgentScope. This category also includes extensions of these tools, for instance, OrcaLoca (yu2025orcaLocallmAgent, ), built on top of Agentless-1.5, improves the issue localization component. We also identify a Code Assistant, Aider, which differs from the previously mentioned assistant in two key aspects: first, it is not implemented as a plug-in but as a command-line tool; second, it is an open-source project developed independently, without backing from a company aiming to commercialize the product or related technologies. The entries associated to Development Assistant achieve the highest median % Resolved within the NCS category with 38.3% on SWE-Bench Lite.
The Under Request (UR) category is associated with 6 entries across both leaderboards. It comprises a variety of product types, including Platform Integration, Code Representation, Code Optimization, and even a Development Plugin. These products are typically available only upon request and are represented by a small number of entries (at most two each). Moreover, submissions in this category are generally limited to a single leaderboard, rather than appearing in both. Notably, the UR category has the highest median % Resolved among all categories.
@@
Finally, we grouped under UN those entries for which no associated product could be identified. This group comprises 16 approaches and 28 entries, with a median % Resolved of 28% on SWE-Bench Lite and 49% on SWE-Bench Verified.
@@
Answer to RQ 1 (Product): Entries associated with Publicly Available Products and Under-Request solutions stand out for their precision, particularly in Verified, where entries associated to Development Assistants reach 62.2% and command-line tools attain 64.6%.

### v2

@@ 3.1.2. Analysis of Submitter Organization
Submission to one or more leaderboards. The Venn diagram in Figure 3 shows which organizations have submitted to SWE-Bench Lite, SWE-Bench Verified, or both. Twenty-two contributors (31%) submitted solutions to both leaderboards, most of them (14) with duplicated submissions (conversely, eight contributors submitted different solutions to both leaderboards). For example, Amazon submitted Amazon Q Developer Agent (v20240430-dev) to both leaderboards. On the other hand, 25 contributors have submitted only to SWE-Bench Lite, and 24 exclusively to SWE-Bench Verified. For instance, IBM has submitted only to SWE-Bench Lite (with two entries), while EPAM has submitted only to SWE-Bench Verified (with five entries).
Analysis of % Resolver rate
Table 2 shows the median and maximum % Resolved achieved by each type of submitter across the two leaderboards.
In SWE-Bench Lite, the highest precision (60.3%) is achieved by ExpeRepair (mu2025experepairdualmemoryenhancedllmbased, ), a submission resulting from a collaboration between academic institutions. A close second (60%) is the submission by a small company, Refact.ai. Notably, recent entries from individual academic groups also surpass the 50% mark, including SWE-Agent (56.67%) and SemAgent-(pabba2025semagentsemanticsawareprogram, ) (51%). Surprisingly, no medium or large company has outperformed these results so far. Across all types of submitters, the median precision remains relatively consistent, ranging from 27% to 29.8%, with the exception of industry submissions, which exhibit a notably higher median (35.8%). We assess whether there are statistically significant differences in precision across submitter types using the Kruskal–Wallis test. The analysis reveals a statistically significant difference between Single Academia and Small Companies (H = 19.89, p = 0.0469). This difference may be influenced by the early submissions to the leaderboard, which were primarily made by academic groups and tended to have lower precision.
In SWE-Bench Verified, the results show a different landscape. Various types of companies —small, medium, and large— submitted results with the highest % Resolved, all exceeding 73%. However, other submitter categories also achieved competitive results. Notably, academia (e.g., Princeton University with SWE-Agent) and the open-source community (e.g., Moatless) reached precision scores above 66%. Excluding the open-source submissions —only two entries, considered outliers—the highest median precision is observed among medium and large companies, driven by submissions from Anthropic, which evaluated its state-of-the-art LLMs featuring hybrid reasoning capabilities. In contrast, the median precision from academia is notably lower than that of the other categories. The Kruskal–Wallis test returned a similar result as SWE-Bench Lite (H = 35.0745, p = 0.0001). Dunn’s post-hoc test further reveals that the precision scores from small, medium and large companies differ significantly from those of academia (pvaluesp\-values 0.00588, 0.00313 and 0.01786, resp.). This finding corroborates the trend discussed above.
@@
Fig 5 shows the leaderboards’ entries arranged according to the date and precision displayed in the leaderboards themselves. Firstly, we observe that the initial entries from both leaderboards come from academia, in particular from the authors of the benchmark. The benchmark was initially used to evaluate the ability of large language models (LLMs) to repair software issues (jimenez2024SWEBenchLLMs, ), and was later applied to assess LLM-based agents (yang2024sweagentagent, ). These early approaches achieved a maximum precision of around 20%. Then, starting in May–June 2024, both leaderboards began receiving submissions from other organizations beyond academia.
In SWE-Bench Lite, as shown in Figure 5(a) we observe that small companies have been constantly pushing the results. Notably, Isoform achieved a state-of-the-art precision of 55% on Jan. 2025, a months later Refact.ai achieved a 60%. Nevertheless, more recently —nearly a year after the leaderboard’s launch— academic submissions have achieved competitive results, led by a collaboration between institutions (mu2025experepairdualmemoryenhancedllmbased, ), and SWE-Agent, from a single academic team (yang2024sweagentagent, ).
In SWE-Bench Verified, as shown in Fig. 5(b), small companies have also consistently driven improvements in precision. For instance, Refact.ai achieved top precision (74.40%) on June 2025. Nevertheless, in 2025, other types of companies reached state-of-the-art performance as well —for example, the medium-sized AI leader Anthropic achieved 73.20%, and the large company Bytedance currently holds the top entry with 75.2%. Submissions from academia have achieved competitive, though not top-ranking, results -driven by Agentless-1.5 (xia2024agentlessdemystifyingllmbasedsoftware, ) (50.8%) and SWE-Agent (yang2024sweagentagent, ) (66%), both developed by single academic teams, as well as by a collaboration between the University of California (UCSB) and Meta, which produced PatchPilot (li2025patchpilot, ) with near state-of-the-art precision of 64.6% on May 2025.
@@
Answer to RQ 1 (Submitters): Entries in the SWE-Bench leaderboard originate from a wide variety of submitters, including (combinations or particularization of) academy, industry and individual developers (who may belong to open source communities). Small companies are dominant, but single academy has almost an equal share of entries in SWE-Bench Lite, although the number of distinct submitters is significantly higher in single academy (with an average of 4 entries per submitter) than in the rest of categories (less than 2 entries per submitter in average). This diversity underscores the benchmark’s accessibility and minimal barriers to participation, demonstrating that it is readily approachable for experimentation by a broad spectrum of contributors. Since the introduction of SWE-Bench Verified, companies (especially small ones) have been driving progress toward state-of-the-art performance.
@@ 3.1.3. Product and Accessibility
PAP Coding Assistant IDE plugin 8 12 33.5 7 15 44.2 10 27 Alibaba Lingma Agent, Amazon Q Developer Agent, AppMap Navie, Augment Agent, Blackbox AI Agent, Bytedance MarsCode Agent, CodeShellAgent, CodeShellTester, Kodu-v1, OpenCSG Starship
Development Assistant Cloud platform 1 1 37 0 0 0 1 1 Patched.Codes Patchwork
@@
Github plugin - Bot 2 2 40.5 1 1 62.2 3 3 AbanteAI MentatBot, CodeStory Aide, CodeStory Midwit Agent
Development Platform Cloud platform 5 6 36.2 7 17 51.8 8 23 Composio SWE-Kit, Emergent E1, Engine Labs, Factory Code Droid, Kortix AI, OpenHands, Solver, devlo
IDE 0 0 0 1 1 71 1 1 Warp
Multiple (Command-line, Plug-in, Cloud) 1 1 60 1 2 72.4 1 3 Refact.ai Agent
Multiple (Plug-in, Cloud) 0 0 0 1 1 70 1 1 Zencoder
Multiple (Plug-in, IDE, Terminal) 0 0 0 1 2 72.9 1 2 TRAE
Other-Inference Service Cloud platform 0 0 0 1 1 40.6 1 1 Nebius AI
Problem Solving Foundation Model 0 0 0 2 6 56.1 2 6 Amazon Nova Premier, Tools + Claude
Total 17 22 36.5 23 47 53 30 69
NCS Agent Framework Command-line tool 1 5 30.7 1 1 70.8 1 6 Moatless Tools
Coding Assistant Command-line tool 1 1 26.3 0 0 0 1 1 Aider
Development Assistant Command-line tool 2 2 38.3 0 0 0 2 2 Lingxi, SuperCoder2
Development Framework Library 0 0 0 1 1 63.4 1 1 AgentScope
Issue Resolution Command-line tool 14 29 27.3 9 23 32.8 17 52 Aegis, Agentless, Agentless + RepoGraph, Agentless Lite, Co-PatcheR, CodeR, DARS Agent, DeepSWE, ExpeRepair, HyperAgent, KGCompass, MCTS-Refine, OrcaLoca + Agentless-1.5, RAG, SWE-Fixer, SWE-RL, SWE-agent
Total 18 37 28.3 11 25 33.6 22 62
B2B Development Platform Command-line tool 0 0 0 1 5 39.6 1 5 EPAM AI/Run Developer Agent
Platform Code Optimization/Improvement Unknown 1 2 24.8 2 3 46.2 2 5 AutoCodeRover, AutoCodeRover-v2
Total 2 3 30.7 3 8 42.9 4 11
UR Development Platform Cloud platform 1 1 30 1 1 52.2 2 2 AIGCode Infant-Coder, Google Jules
Platform Code Optimization/Improvement Cloud platform 0 0 0 1 1 32 1 1 Artemis Agent
Platform Code Representation None 0 0 0 1 1 53.2 1 1 Bracket.sh
Platform Integration None 1 2 45 0 0 0 1 2 Isoform
@@
None None 13 14 35.7 10 16 46.1 19 30 Aime-coder v1, Bytedance AutoSE, CORTEXA, Codart AI, CodeFuse-AAIS, Codev, Gru, Honeycomb, IBM AI Agent SWE-1, IBM Research Agent-101, Learn-by-interact, MASAI, PatchKitty_PatchPilot, SIMA, SemAgent, Skywork-SWE, nFactorial, reproducedRG, ugaiforge
Total 13 14 35.7 10 16 46.1 19 30
@@
Table 3 presents the number of approaches (#A) and entries (#E) per product availability category, along with their median precision (% Resolved). We analyse below the contents of each category.
@@
Publicly Available Product (PAP). It includes publicly accessible commercial tools and accounts for 69 entries across both leaderboards, representing 30 distinct approaches. This category includes products with different purposes, namely Coding Assistants, Development Platforms and Development Assistants. The most common product type is Coding Assistant, comprising 10 products available as IDE plug-ins and linked 27 entries. These include, for example, Amazon Q Developer, Alibaba Lingma, and AppMap Navie, which are evaluated on SWE-Bench and available as extensions for popular development environments such as IntelliJ JetBrains IDE and Visual Studio Code. The second major product type is Development Platform, typically delivered as cloud-based services. These platforms generally allow developers to specify new features using natural language and integrate directly with code repositories (e.g., GitHub). There are 23 entries linked to these platforms across both leaderboards and include solutions such as OpenHands, Solver, and Factory Code Droid. Other solutions such as Refact.ai, TRAE or Zencoder offer different product forms in addition to the cloud Services, such as IDE plugins. Notably, the submissions related Development Platform products achieve the highest median % Resolved across both leaderboards. A third type includes Development Assistant, which helps developers in particular development task. Three of the entries, such as CodeStory Aide (later rebranded as AgentFarm), materialize this assistant as automated bots integrated into GitHub workflows, interacting with developers via pull requests, while the two others as a command-line tool and a cloud platform. The submissions related PAP products achieve the highest median % Resolved across both leaderboards (36.5% and 53%), closely followed by those from UR (Under Request) category.
@@
The Non-Commercial Solutions (NCS) category accounts for 62 entries across both leaderboards, corresponding to 22 distinct approaches. The dominant product type is Issue Resolution, comprising 17 approaches such as Agentless, SWE-agent, HyperAgent, and AgentScope. This category also includes extensions of these tools, for instance, OrcaLoca (yu2025orcaLocallmAgent, ), built on top of Agentless-1.5, improves the issue localization component. We also identify a Code Assistant, Aider, which differs from the previously mentioned assistant in two key aspects: first, it is not implemented as a plug-in but as a command-line tool; second, it is an open-source project developed independently, without backing from a company aiming to commercialize the product or related technologies.
@@
The Under Request (UR) category is associated with 6 entries from 5 approaches across both leaderboards. It comprises a variety of product types, including Platform Integration, Code Representation, Code Optimization, and even a Development Plugin. These products are typically available only upon request and are represented by a small number of entries (at most two each). Moreover, submissions in this category are generally limited to a single leaderboard, rather than appearing in both.
Finally, we grouped under UN (unknown) those entries for which no associated product could be identified. This group comprises 16 approaches and 28 entries, with a median % Resolved of 28% on SWE-Bench Lite and 49% on SWE-Bench Verified.
@@
Temporal Evolution of % Resolved
Fig. 6 presents the leaderboard entries over time, showing their submission dates and the corresponding precision scores. In SWE-Bench Verified, entries associated with publicly available products (PAP) have consistently achieved state-of-the-art results up to the present date. In contrast, the trend in SWE-Bench Lite differs: since early 2025, only a few PAP-related entries have been published, with the majority originating from non-commercial systems (NCS). This may be attributed to the observation that industrial actors —who are more likely to release commercial products— tend to submit more to SWE-Bench Verified than to SWE-Bench Lite, as previously discussed.
Answer to RQ 1 (Product): Entries in the SWE-Bench leaderboard are made available to the community in different ways: publicly accessible commercial tools, non-commercial solutions, available under request or as B2B solutions. Entries associated with Publicly Available Products and Under-Request solutions stand out for their precision, particularly in SWE-Bench Verified. A non-negligible number of entries (19% over total) couldn’t be classified.
(a) SWE-Bench Lite
(b) SWE-Bench Verified
Figure 6. Evolution of % Resolved per product availability. The availabilities are PAP: Publicly Available Product, UR: Upon Request, B2B: Business-to-Business, NCS: Non-Commercial Solution, UN: Unavailable.